# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models

### v4

@@ 3.1. Code Completion Task
3.2. Target-specific Attack
@@
(1) $\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\}.\end{split}$
3.3. Threat Model
@@
Attack Scenarios. We consider common scenarios where users introduce external code by copying or installing packages. The attacker generates a concise non-functional perturbation containing malicious instructions and inserts it into the victim’s code context via the external code. When the user requests the Code LLM to perform code completion based on code context, the perturbation will induce the model to generate an unsafe output at the specified location. In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by existing main-stream detection methods (e.g., static analysis (Li et al., 2018b)) since they directly insert malicious functional code, to implant malicious behaviors, our attack scenario is more threatening because the attacker can circumvent detection methods by introducing only non-functional malicious code. Besides introducing code to the victim’s local project, attackers can upload their malicious payload codes to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020). Code in the knowledge base can also appear in the context of the Code LLM’s prompts.
Attacker’s Capabilities. We assume that the attacker can obtain the parameters of Code LLM to generate malicious perturbations. Arguably, when developers build integrated applications using open-source models (such as CodeGemma (Team et al., 2024), CodeLlama (Roziere et al., 2023), or CodeGeeX (Zheng et al., 2023)), obtaining the corresponding model parameters is feasible. In this scenario, TPIA can achieve optimal performance. In this paper, we also consider gray-box (i.e., with unknown model version and system prompts) and black-box (i.e., with closed-source commercial models) scenarios, based on two widely-used commercial Code LLM-integrated applications in Section 5.3. We conduct successful attacks against these applications via perturbations generated with a local open-source Code LLM, confirming the practical threat of our TPIA in the real world.
On the other hand, the attacker needs to submit the payload code into the code repository and prevent it from being detected. Since our threat model uses non-functional code to achieve objectives, the payload code does not need to be designed with malicious functionality. In other words, the attacker can freely design the functional part of the payload code to ensure successful submission without worrying about detection by methods based on malicious functionality. In contrast, the perturbation part of the payload needs to be stealthy enough to avoid detection by victim users.
@@
Attacker’s Goals. The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code). This tuple follows the definition of target-specific attacks in Section 3.2. Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$. The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$. To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$. It can be formulated as below:
(2) $\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$
@@
Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section 3.2 and the generated perturbation $\bm{P}$ should satisfy Equation (1).
@@
Figure 2. The overall pipeline of our TPIA attack. It is executed with two attack steps. (1) Objective Transfer. The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective. (2) Perturbation Injection. The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub (Git, 2024) and StackOverflow (Sta, 2024)) and introduced to the victim’s code context via various paths, such as direct copy and installing packages. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
4. Methodology
4.1. Overview
We introduce our target-specific prompt injection attack (TPIA). In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2. Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section 4.2). In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section 4.3). The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section 4.4); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section 4.5). We describe their technical details in the following parts.
4.2. Task Construction
@@
Property 2: Accurate Code LLMs’ Response. The response of the Code LLMs should be the target position. As stated in Section 3, the target code should be inserted after the target position to achieve the malicious objective. Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
@@
To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”. To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”. The other three task components are automatically generated with the following stages. Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code. Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”. In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”. Finally, TPIA constructs the code unrelated to the task, such as “generate_random_string”, and uses it as the noise-like context code. Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
Figure 3. The example of task construction that induces Code LLM to use encryption mode with low security. In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
4.3. Malicious Perturbation Generation
In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code. Utilizing the code completion task constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer. Through a greedy search approach, the optimal perturbation is derived. The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
@@ 4.3. Malicious Perturbation Generation
Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation. Considering that our enhancement method adds static tokens to the perturbation (Section 4.5), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$. After determining the input and output of the code completion task, TPIA performs continuous iterations to optimize the perturbation. Specifically, since the attack is ultimately executed by the perturbation embedded in the external source code, in each iteration, TPIA only replaces the adversarial part of the perturbation token by token (Line 4). We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
(3) $\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$
@@
(4) $\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$
@@
To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
(5) $e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$
@@
where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss. TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7). This has been proven to be highly effective in existing NLP tasks (Zou et al., 2023; Wallace et al., 2019). Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9). It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
@@
In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger (Wallace et al., 2019) and GCG (Zou et al., 2023). We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection). However, its effect is still very limited due to the difficulties of targeted malicious tasks. For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%. To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA. More details about the comparison between existing methods and TPIA are discussed in Appendix A.1.
@@
Figure 4. Loss function distortion during the forward reasoning process. When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence. However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective.
4.4. Forward Reasoning Enhancement
@@
This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. Specifically, the calculation of the loss function in Malicious Perturbation Generation will encounter the distortion issue during the actual reasoning process of the Code LLMs. As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context. In our experiment, a malicious perturbation might have a much smaller loss value compared to a successful manually-crafted perturbation, yet the former fails to produce the desired output. This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions. This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
@@
To alleviate this problem, we propose a forward reasoning enhancement technique. Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning. Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error. Its formulation is shown as follows:
(6) $\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$
@@
Figure 5. Examples of the keyword-based perturbation and other perturbations. The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage. Our keyword-based perturbation can balance both effectiveness and stealthiness.
4.5. Keyword-based Perturbation Design
@@
This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack. Specifically, since the content of the perturbation and the payload code is not logically consistent, the attacker must enhance perturbation’s stealthiness as much as possible to avoid human detection. From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack. Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent. However, although we can use some optimization methods to obtain the optimal perturbation, it is still difficult to use illegible tokens to express complex instructions. A typical failure case is that even after the optimization, the attack objective cannot be completely transferred into the perturbation, resulting in attack invalidation. A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues. First, the malicious perturbation cannot gain any advantage in terms of the token count. Second, this leads to ineffective optimization efforts. As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds. Additionally, the complex tokens generated after adversarial computation often occupy more vision space.
To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design. This method allows the malicious perturbation to consume fewer tokens to convey malicious instruction, enhancing the attack stealthiness. Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens. Specifically, we need to use non-sensitive information that does not directly contain malicious features, which we call keywords, as static tokens for perturbation. For instance, as shown in Figure 5, part of the target function name (TR) is the key information, which is not directly related to the malicious objective (data deletion). The tokens of keywords are selected from the position code and target code after each perturbation optimization process, and the content is ensured to be non-malicious by limiting the number and position of selected tokens. The details of this method is discussed in Appendix A.2. By initializing the perturbation as “keywords<AT><AT><AT><AT>”, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
5. Evaluation

### v5

@@ III-A Code Completion Task
III-B Target-specific Attack
@@
$\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\},\end{split}$ (1)
where $\oplus$ means contacting the two token sequences.
Figure 2: The overall pipeline of our TPIA attack. It is executed with two attack steps. (1) Objective Transfer. The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective. (2) Perturbation Injection. The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub [3] and StackOverflow [5]) and introduced to the victim’s code context via various paths, such as direct copy and installing packages. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
III-C Threat Model
@@
Attack Scenarios. We consider common scenarios where users introduce external code by copying or installing pack- ages [22, 41]. The attacker generates a non-functional perturbation to circumvent detection methods (e.g., static analysis [33] and sandbox [9]) and inserts it into victim’s code context via the external code. When the user requests the Code LLM to perform code completion, the perturbation will induce the model to generate an unsafe output at the specified location. Besides introducing code to the victim’s local project, attackers can also upload their codes containing perturbation to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27]. Perturbation in the knowledge base can also appear in the context of the Code LLM’s prompts.
Attacker’s Capabilities. We assume that the attacker can obtain the parameters of an assistive Code LLM to generate malicious perturbations. This assumption is consistent with many adversarial methods [63, 50, 24]. Specifically, the attacker ideally obtains the open-source model corresponding to the target application, or uses the most effective assistive model in gray-box and black-box scenarios. On the other hand, perturbations can be detected and removed. Therefore, the attacker needs to employ non-functional perturbations rather than the functional code commonly used in supply chain attacks [41], since the latter can be easily detected by existing methods (e.g., static analysis [33] and sandboxing [9]).
@@
Attacker’s Goals. The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code). This tuple follows the definition of target-specific attacks in Section III-B. Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$. The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$. To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$. It can be formulated as below:
$\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$ (2)
@@
Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section III-B and the generated perturbation $\bm{P}$ should satisfy Equation (1).
@@
IV Methodology
IV-A Overview
We hereby briefly introduce our target-specific prompt injection attack (TPIA). In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2. Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section IV-B). In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section IV-C). The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section IV-D); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section IV-E). We describe their technical details in the following parts.
IV-B Task Construction
@@
Property 2: Accurate Code LLMs’ Response. The response of the Code LLMs should be the target position. As stated in Section III, the target code should be inserted after the target position to achieve the malicious objective. Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
@@
To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”. To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”. The other three task components are automatically generated with the following stages. Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code. Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”. In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”. Finally, TPIA constructs the code unrelated to the task and uses it as the noise-like context code. Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
Figure 3: The example of task construction that induces Code LLM to use encryption mode with low security. In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
IV-C Malicious Perturbation Generation
In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code. Utilizing the task code constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer. Through a greedy search approach, the optimal perturbation is derived. The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
@@ IV-C Malicious Perturbation Generation
Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation. Considering that our enhancement method adds static tokens to the perturbation (Section IV-E), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$. After determining the input and output of the code completion task, TPIA starts to optimize the $x_{\text{adv}_{1:l}}$ token by token in each iter (Line 4). We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
$\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (3)
@@
$\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$ (4)
@@
To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set of the Code LLM $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$ (5)
@@
where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss. TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7). This has been proven to be highly effective in existing NLP tasks [63, 50]. Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9). It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
@@
In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger [50] and GCG [63]. We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection). However, its effect is still very limited due to the difficulties of targeted malicious tasks. For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%. To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA.
@@
Figure 4: Loss function distortion during the forward reasoning process. When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence. However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective in practical scenarios.
IV-D Forward Reasoning Enhancement
@@
This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context. This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions. This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
@@
To alleviate this problem, we propose a forward reasoning enhancement technique. Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning. Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error. Its specific formulation is shown as follows:
$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (6)
@@
Figure 5: Examples of the keyword-based perturbation and other perturbations. The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage. Our keyword-based perturbation can balance both effectiveness and stealthiness.
IV-E Keyword-based Perturbation Design
@@
This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack. From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack. Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent. However, accurately conveying instructions using illegible tokens remains a significant challenge. A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues. First, the malicious perturbation cannot gain any advantage in terms of the token count. Second, this leads to ineffective optimization efforts. As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds.
To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design. Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens. Specifically, before each round of perturbation optimization in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords. With the optimization step $s$, the equation can be formulated as follows:
$\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$ (7)
where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6. To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{\emph{target code}},\text{\emph{position code}}\}$. As shown in Figure 5, part of the target function name (TR) is the keywords, which is not directly related to the malicious objective (data deletion). By using a small number of statistical tokens, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
V Experiments